A rental company charges $30 a day and 15 cents per mile for renting a car. Let y be the total charges (in dollars) for a car for one day and x be the miles driven. The equation for the relationship between x and y is:
Y = 30 + .15X

a. How much will a person pay who rents a car one day and drives it 100 miles?
b. Suppose each of 20 persons rents a car from this agency for one day and drives it 100 miles. Will each of them pay the same amount for renting a car for a day or is each person expected to pay a different amount? Explain.
c. Is the relationship between X and Y exact or nonexact?
d. Identify the dependent and independent variables.

Answer by stanbon(75887)

A rental company charges $30 a day and 15 cents per mile for renting a car.
Let y be the total charges (in dollars) for a car for one day and x be the miles driven. The equation for the relationship between x and y is:
Y = 30 + .15X
a. How much will a person pay who rents a car one day and drives it 100 miles?
y = 30+0.15*100 = 30 + 15 = $45
-------------------------------------
b. Suppose each of 20 persons rents a car from this agency for one day and drives it 100 miles. Will each of them pay the same amount for renting a car for a day or is each person expected to pay a different amount? Explain.
Each one pays the same as the others.
-------------------------------------
c. Is the relationship between X and Y exact or nonexact?
exact
---------------
d. Identify the dependent and independent variables.
independent: milage
dependent: cost
============================
Cheers,
Stan H.